# Multi-Agent Diagnostics for Robustness via Illuminated Diversity

### v1

@@ Appendix B Environment Details
We use the enhanced observation space as described in TiZero (Lin et al., 2023), consisting of $268$-dimensional vector including information.
@@ Appendix C Implementation Details
The policy network is made up of six different multi-layer perceptrons (MLPs), each having two fully-connected layers, including one specifically for the ’player ID’, to encode every part of the observation individually. The MLP layers have a hidden size of 64. The hidden features extracted are brought together and then handled by an LSTM layer to give the agent memory, with the hidden size for this layer being 256. Every hidden layer is equipped with layer normalization and ReLU non-linearities. The orthogonal matrix is used for initializing parameters, and the learning process is optimized with the Adam optimizer. Similar to the original implementation, illegal actions are masked out by making their selection probability zero. The action output layer utilizes a softmax layer and is formed with a 19-dimension vector.

### v2

@@ Appendix B Environment Details
We use the enhanced observation space as described in TiZero (Lin et al., 2023), consisting of a $268$-dimensional vector including information.
@@ Appendix C Implementation Details
The policy network is made up of six different multi-layer perceptrons (MLPs), each having two fully connected layers, including one specifically for the ’player ID’, to encode every part of the observation individually. The MLP layers have a hidden size of 64. The hidden features extracted are brought together and then handled by an LSTM layer to give the agent memory, with the hidden size for this layer being 256. Every hidden layer is equipped with layer normalisation and ReLU non-linearities. The orthogonal matrix is used for initialising parameters, and the learning process is optimized with the Adam optimizer. Similar to the original implementation, illegal actions are masked out by making their selection probability zero. The action output layer utilises a softmax layer and is formed with a 19-dimension vector.